3. What does this code do? Write an explanation in English python n=1 while (n <1000): print(n) n-n
Answer #1

#! is called shebang, is used mostly in Unix based systems. It is followed by the full path to the interpreter such as /bin/bash or /usr/bin/python. All scripts under Linux execute using the interpreter specified on a first line.

n=1 :- initial/ start value of n.

while(n<1000):     #run this loop till n is smaller than 1000, If n becomes greater or equal to 1000,

   # this loop will be over and our statement will go pass this loop

    print(n)           # printing the value of n in each iteration

    n = n * 5         #or it can be written like n *= 5 . No spacing between asterisk and assignment operator

                             #we are multiplying n to 5 and assigning the value to the n itself.

                              #Main thing to notice is that n is multiplied by 5 and assign to n ,

                              #after first iteration inside the loop n has become 5. but still less than 1000. Till

                              #that time n will be updated.

Here:

Before Entering the loop n=1

(1 < 1000 ) #here condition is checked, if condition is True, then we go into the loop else we skip the loop

After the entering for the first time(iteration 1): n = n*5 i.e, n=1*5=5. operations are done from right handside of the assignment operator

(5 < 1000)

iter2: n = n*5 = 5*5=25

(25<1000)

iter3: n=n*5 = 25 *5=125

(125 < 1000)

iter4: n=n*5 = 125 * 5 = 625

(625<1000)

iter5: n= n*5 = 625 * 5 = 3125

(3125 < 1000) ###FALSE CONDITION, we break out of the loop.

Hope You get your answer.